OpenStudy (anonymous):

acb be similar to dce. ac=32,ad=8 and ce=18 what is the length of be?

OpenStudy (anonymous):

this are triangles

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

then use property of similarity of two triangles

OpenStudy (anonymous):

I don't know how to.

OpenStudy (anonymous):

i will tell you okay do what i am saying okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if two triangles are said to be similar then their corresponding sides are prapotional

OpenStudy (anonymous):

can you draw diagram of question

OpenStudy (anonymous):



OpenStudy (anonymous):

you can get value of dc by subtracting ac - ad

OpenStudy (anonymous):

means cd is 32 - 8= 24 okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

according to property we can say ad/cd = cb / ce

OpenStudy (anonymous):

okay but where does be come into play?

OpenStudy (anonymous):

i think you first understand property of similarity okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

triangle abc is similar to triangle cde okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so both corresponding sides are praportional

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that is why ad/cd = cb / ce

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then solve on putting values

OpenStudy (anonymous):

would it be 0.01851852?

OpenStudy (anonymous):

32/24=cd / 18

OpenStudy (anonymous):

cd = 32 * 24 / 18

OpenStudy (anonymous):

42.6666667?

OpenStudy (anonymous):

to find be you subtract bc - ce

OpenStudy (anonymous):

means 46.66 - 18

OpenStudy (anonymous):

28.66

OpenStudy (anonymous):

this is your ans

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

your welcome